1. How many sheets of paper, weighing 2 grams each, can be put in an envelope weighing 4 grams if the total weight must not exceed 29 grams?
Answer by mananth(16949)

let number of sheets be x
weight of each sheeet = 2gm.
enevelope = 4 g
..
total weight =29
2x+4 =29
-4
2x+4-4 =29-4
2x=25
/2
x= 25/2
= 12.5
12 sheets can be out inside